# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ Abstract
The valley degree of freedom in honeycomb crystals such as 2D semiconductors, graphene and bilayer graphene is a promising carrier of quantum information alongside spin and charge. This is all the more true since gate-controlled single-particle quantum dots (QDs) have been demonstrated in bilayer graphene (BLG), paving the way for the realisation of spin and valley qubits. Although long spin relaxation times have recently been reported in BLG QDs, nothing is known about single-particle valley lifetimes. Here we report single-particle valley relaxation times ($T_{1}$ times) exceeding several microseconds in electrostatically defined BLG QDs. The observed dependence of $T_{1}$ on the perpendicular magnetic field can be understood qualitatively and quantitatively by a model in which $T_{1}$ is limited by electron-phonon coupling. We identify coupling to acoustic phonons via the bond length change and via the deformation potential as the limiting mechanisms.
@@
Bernal-stacked BLG comes as a gapless semimetal, where electrons and holes can be described as massive Dirac fermions Novoselov et al. (2006). However, when applying an out-of-plane electric displacement field $D$, the inversion symmetry of the crystal lattice is broken, as the on-site energy of carbon atoms of the top layer becomes different from that of the atoms of the bottom layer (see Fig. 1a) McCann and Koshino (2013). This leads to the opening of a band gap at the two valleys, $K$ and $K^{\prime}$ (see Fig. 1b), which depends on the strength of the symmetry breaking potential, i.e., on $D$  Oostinga et al. (2008); Zhang et al. (2009); Icking et al. (2022), resulting in a highly tunable band structure that allows for gate-defined charge carrier confinement Eich et al. (2018); Banszerus et al. (2020a, 2018). The broken inversion symmetry also leads to a finite Berry curvature, $\Omega$, near the K points, where $\Omega$ has opposite signs at the $K$ and $K^{\prime}$ points and has mirror symmetry for electrons and holes McCann and Koshino (2013); Banszerus et al. (2023) (see Fig. 1c). The Berry curvature gives rise to a valley-dependent anomalous velocity term leading to the valley-Hall effect in extended BLG samples Shimazaki et al. (2015); Sui et al. (2015) and to finite out-of-plane magnetic moments in BLG QDs. These topological orbital magnetic moments, which have opposite signs for $K$ and $K^{\prime}$, couple to an external out-of-plane magnetic field and are the origin of the valley Zeeman effect in BLG QDs Banszerus et al. (2021a); Tong et al. (2021).
@@
Figure 1: Electrostatically defined single-particle QD in BLG. a Lattice structure of bilayer graphene highlighting the top and bottom layer and a symmetry breaking displacement field, $D$. The electrostatic confinement potential $U$ (shown in the lower panel) allows to form a QD. b Band structure of gapped BLG. Close to the $K$ (blue/red) and the $K^{\prime}$ (green/yellow) valley a band gap opens. c The broken inversion symmetry results in a finite Berry curvature, $\Omega$, near the $K$ and $K^{\prime}$ points, which has opposite sign for the two valleys and for electrons and holes. d Single-particle spectrum of a BLG QD. At $B_{\perp}=0$, the four states are grouped into two spin and valley unpolarized Kramers’ doublets split by the Kane Mele SO gap, $\Delta_{\text{SO}}$. A finite $B_{\perp}$ couples to both the spin and valley magnetic moments, resulting in a spin and valley Zeeman effect, leading to additional energy splittings of $\Delta E_{\text{s}}=g_{\mathrm{s}}\mu_{\mathrm{B}}B$ and $\Delta E_{\text{v}}=g_{\mathrm{v}}\mu_{\mathrm{B}}B$. The arrows depict the transition energies between the ground state |$\ket{K^{\prime}\uparrow}$ and the three excited states.
@@
To successfully create gate-defined QDs in BLG, the electronic wave function needs to be confined by a potential $U(\textbf{r})$ in real space (see Fig. 1a) and will be distributed near the $K$ and $K^{\prime}$ points, in k-space. A single-electron or single-hole QD can then be described by the Hamiltonian $H_{\rm QD}=H_{\rm BLG}+H_{\rm Z}+H_{\rm SO}+U({\bf r})$ Wang and Burkard . Here, $H_{\rm BLG}$ represents the effective $4\times 4$ Hamiltonian of bulk BLG near the $K$- and $K^{\prime}$-points based on the sublattice and layer degrees of freedom and includes the bulk valley Zeeman effect (see Methods), which will be further modified by the confinement $U({\bf r})$. $H_{\rm Z}$ denotes the spin Zeeman coupling. $H_{\rm SO}$ describes the intrinsic Kane-Mele spin-orbit (SO) coupling, which lifts the zero $B$-field degeneracy of the four single-particle states, leading to the formation of two Kramers’ pairs with opposite spin and valley quantum numbers, (|$\ket{K\uparrow}$,|$\ket{K^{\prime}\downarrow}$) and (|$\ket{K^{\prime}\uparrow}$, |$\ket{K\downarrow}$), separated by the SO gap, $\Delta_{\mathrm{SO}}$. In BLG quantum devices $\Delta_{\mathrm{SO}}$ has typically values in the range of $40-80~{}\,\mu$eV Banszerus et al. (2020a, 2021a, 2023); Kurzmann et al. (2021); Tong et al. (2022). In Fig. 1d we show the single-particle spectrum of a BLG hole QD as function of the out-of-plane magnetic field, $B_{\perp}$. As $B_{\perp}$ couples to both the spin and the valley magnetic moments, we observe linear energy shifts given by $E(B_{\perp})-E(0)=(\pm g_{\mathrm{s}}\pm g_{\mathrm{v}})\mu_{\mathrm{B}}B_{\perp}/2$ Knothe and Fal’ko (2018). Here, $\mu_{\text{B}}$ is the Bohr magneton, $g_{\mathrm{s}}\approx 2$ is the spin g-factor and the valley g-factor, $g_{\mathrm{v}}$, quantifies the strength of the Berry-curvature induced valley magnetic moment, which can be tuned by the confinement potential of the QD in a range typically between $g_{\text{v}}\approx 10$ and 70 Tong et al. (2021); Möller et al. (2021). All this makes the valley degree of freedom well accessible and allows significant valley polarisation at reasonably small $B_{\perp}$-fields (see separation of |$\ket{K^{\prime}}$ and excited |$\ket{K}$ states in Fig. 1d).
@@
Figure 2: BLG quantum dot device and measurement of the single-particle spectrum. a False-colored scanning electron microscopy image of the gate structure of the device. A channel is formed between source (S) and drain (D) using two split gates (SGs). The potential along the channel is controlled using two layers of finger gates (FGs). One FG (red) is used as a plunger gate (PG) for the QD, while the neighboring FGs (yellow and blue) are used to tune the tunneling barriers between the QD and the reservoirs. The PG is connected to a bias tee for applying AC and DC signals to the same gate. b Current through the device as function of $V_{\text{PG}}$ at a source-drain bias voltage of $V_{\mathrm{SD}}=200~{}\mu$V. The n-type channel is pinched off close to $V_{\text{PG}}=-6.8V$. Upon further decreasing $V_{\text{PG}}$, a hole QD is formed, and Coulomb resonances appear when additional holes are added to the QD (see labels). c Top: Schematic of the square pulse applied to the PG which is characterized by the voltages $V_{\text{i}}$ and $V_{\text{m}}$ and the times $\tau_{\text{i}}$ and $\tau_{\text{m}}$. Bottom: Schematic of the QD states relative to the electrochemical potentials of the leads. d Excited state (ES) spectroscopy using transient current measurements. The average number of charge carriers $\langle n\rangle$ tunneling through the QD per pulse is plotted as function of $\Delta V_{\mathrm{PG}}$ for different $V_{\mathrm{AC}}$. Traces are offset for clarity. Current via the ground state |$\ket{K^{\prime}\uparrow}$ occurs at $\Delta V_{\mathrm{PG}}=0~{}$mV. Orange and yellow arrows and dashed lines highlight transient currents via excited states (|$\ket{K^{\prime}\downarrow}$, |$\ket{K\uparrow}$ and |$\ket{K\downarrow}$). e Energy $\Delta E$ of the ES relative to the GS as a function of $B_{\perp}$. From a fit to the energy of the spin ES (yellow), $\Delta_{\text{SO}}$ and $g_{\mathrm{s}}$ is extracted, while a fit according to $\Delta_{\text{SO}}/2+(g_{\mathrm{v}}+g_{\mathrm{s}}/2)\mu_{\mathrm{B}}B$ (orange dashed line) yields $g_{\mathrm{v}}$. The solid lines indicate the energies of the valley ESs deduced from the fits. The inset shows a close-up of the low-$B_{\perp}$ regime.
@@
The fabricated device used to confine single charge carriers in a BLG QD is depicted by the scanning electron micrograph in Fig. 2a. The QD device consists of a flake of BLG encapsulated by two crystals of hexagonal boron nitride (hBN) and placed on a graphite flake acting as a back gate. On top of the van-der-Waals heterostructure, metallic split gates (SGs) are used to gap out the BLG underneath the SG areas, resulting in a narrow n-type conductive channel connecting the source and drain leads (see contacts in Fig. 2a). To confine charge carriers, the band edge profile along the channel can be adjusted using two layers of interdigitated finger gates (FGs) Banszerus et al. (2018); Eich et al. (2018). One of the FGs is used as a plunger gate (PG) to tune the QD (see red FG in Fig. 2a) and allows to locally overcompensate the channel potential set by the back gate. The width of the PG measures about 70 nm and the separation of the SGs is around 80 nm, which set an upper limit of the QD radius $r$ to around 30 to 40 nm. The DC potential applied to the plunger gate, $V_{\text{PG}}$, allows to control the charge carrier occupation of the QD down to the last hole (see Fig. 2b). Additionally, to study transient transport through the QD, an AC potential, $V_{\mathrm{AC}}$, can be applied to the PG via a bias tee (see Fig. 2a). To maximize the transient currents and to study the relaxation dynamics of the QD states, the FGs adjacent to the PG are used to reduce the tunnel coupling between the QD and the left and right reservoir, $\Gamma_{\mathrm{L}}$ and $\Gamma_{\mathrm{R}}$, respectively (see colored FGs in Fig. 2a) Hanson et al. (2003); Banszerus et al. (2021b, 2022); Gächter et al. (2022).
@@
To study the relaxation dynamics of an excited valley state, we first investigate the single-particle spectrum of the QD. For that purpose, we perform excited state transient current spectroscopy measurements by applying a square pulse with a frequency, $f$, (duty cycle 50%) to the PG (see Fig. 2c and Methods for details). Fig. 2d shows the average number of charge carriers that are tunneling through the QD per pulse cycle, $\langle n\rangle/\text{pulse}=I/(fe)$, with the elementary charge $e$, as a function of the pulse amplitude, $V_{\mathrm{AC}}$, and the DC plunger gate voltage, $\Delta V_{\mathrm{PG}}$, relative to the Coulomb peak position at $V_{\mathrm{AC}}=0$. The pulse frequency has been fixed at $f=5~{}$MHz and the magnetic field at $B_{\perp}=300$ mT. At finite $V_{\mathrm{AC}}$, transport via the ground state (GS) may occur when the GS resides within the bias window during either part of the square pulse ($\tau_{\mathrm{i}},\tau_{\mathrm{m}}$), resulting in a splitting of the GS Coulomb peak (|$\ket{K\uparrow}_{i}$ and |$\ket{K\uparrow}_{m}$ujisawa et al. (2001); Banszerus et al. (2021b). Once $V_{\mathrm{AC}}$ becomes large enough such that an excited state (ES) enters the bias window, a transient current via the ES contributes to the overall current and shows up as a resonance in Fig. 2d (see coloured dashed lines). From the positions of the two prominent ES resonances in Fig. 2d, we can extract their corresponding ES energies (see Methods) and measurements as shown in Fig. 2d have been recorded for various $B_{\perp}$-fields. Fig. 2e depicts the energy difference between the ground state (|$\ket{K^{\prime}\uparrow}$) and the first spin ES (yellow data points, |$\ket{K^{\prime}\downarrow}$) and the valley ESs, (red data points |$\ket{K\downarrow}$, |$\ket{K\uparrow}$) as a function of $B_{\perp}$. The energy splitting of the spin ES and the GS increases linearly with $B_{\perp}$ due to the spin Zeeman effect. A fit to the data yields a spin g-factor of $g_{\text{s}}=2.0\pm 0.2$ and a zero-field splitting of $\Delta_{\text{SO}}=75~{}\mu$eV in agreement with the slightly proximity-enhanced Kane-Mele SO coupling Banszerus et al. (2021a, 2023); Kurzmann et al. (2021). Due to the finite peak width, the energy of the two valley ESs cannot be determined independently. Thus, the data has been fit considering the average energy splitting with a slope corresponding to $g_{\mathrm{v}}+g_{\mathrm{s}}/2$. A valley g-factor of $g_{\mathrm{v}}=30.2\pm 0.2$ has been determined, similar to values reported in earlier works Banszerus et al. (2021a); Tong et al. (2021).
@@
Figure 3: Measurement of the valley relaxation time. a Top: Schematic of the three-level pulse scheme applied to the PG which is characterized by the voltages $V_{\text{i}},V_{\text{h}}$ and $V_{\text{m}}$ and the times $\tau_{\text{i}},\tau_{\text{h}}$ and $\tau_{\text{m}}$. Bottom: Schematic of the QD states relative to the electrochemical potentials in the leads (see text for details). b Current $I$ as a function of $V_{\text{PG}}$ while the pulse sequence in a is applied. The valley $T_{1}$ time is derived from the amplitude of |$\ket{K\uparrow}_{\text{m}}$. Dashed curves are Lorentzian fits to the peaks. The measurement has been taken at $B_{\perp}=0.22$ T and $V_{\mathrm{SD}}=10~{}\mu$V. c Relative occupation probability of |$\ket{K\uparrow}$ after the holding pulse, $P_{\ket{K\uparrow}}(\tau_{\text{h}})/P_{\ket{K\uparrow}}(0)$ as a function of the holding time, $\tau_{\text{h}}$. Data has been taken at $B_{\perp}=$ 0.175, 0.25 and 0.45 T, respectively. The traces are offset for clarity.
@@
Next, we investigate the relaxation dynamics of the observed single-particle valley ES. For that purpose, we apply a three-level pulse scheme to the PG and measure the tunneling current through the QD Banszerus et al. (2022). The pulse scheme, depicted in Fig. 3a, is characterized by the lengths of each pulse step ($\tau_{\mathrm{i}}$, $\tau_{\mathrm{h}}$ and $\tau_{\mathrm{m}}$) and the corresponding voltages ($V_{\mathrm{i}}$, $V_{\mathrm{h}}$ and $V_{\mathrm{m}}$). During $\tau_{\mathrm{i}}$, the QD is initialized in the empty state. Subsequently, during $\tau_{\mathrm{h}}$, the GS |$\ket{K^{\prime}\uparrow}$, the spin ES |$\ket{K^{\prime}\downarrow}$ and the two valley ESs |$\ket{K\uparrow}$ and |$\ket{K\downarrow}$ are tuned below the electrochemical potentials of the source and drain leads. After the characteristic tunneling time on the order of $1/(\Gamma_{\mathrm{L}}+\Gamma_{\mathrm{R}})$ either the GS or one of the three ESs will be occupied by a single charge carrier. A charge carrier in an ES has the chance to relax into an energetically lower lying state by either spin or valley relaxation with a characteristic relaxation time, $T_{1}$. Finally, during $\tau_{\mathrm{m}}$, we perform a valley selective readout measured by aligning the |$\ket{K}$ states in the bias window. Only charge carriers occupying one of the two |$\ket{K}$ states, which have not relaxed into a |$\ket{K^{\prime}}$, can tunnel out and contribute to the transient current.
Fig. 3b shows the current, $I$, through the QD as a function of $V_{\mathrm{PG}}$ while applying the pulse sequence depicted in Fig. 3a. The three peaks labeled |$\ket{K^{\prime}\uparrow}_{\mathrm{i}}$, |$\ket{K^{\prime}\uparrow}_{\mathrm{h}}$ and |$\ket{K^{\prime}\uparrow}_{\mathrm{m}}$ correspond to GS transport during each of the three pulse steps. Furthermore, transient currents via the three ESs, |$\ket{K\uparrow}_{\mathrm{m}}$, |$\ket{K\downarrow}_{\mathrm{m}}$ and |$\ket{K^{\prime}\downarrow}_{\mathrm{m}}$, can be observed during $\tau_{\mathrm{m}}$. The relaxation time $T_{1}$ of |$\ket{K\uparrow}$ into an energetically lower lying state can be probed by measuring the amplitude of the |$\ket{K\uparrow}_{\mathrm{m}}$ peak as a function of the holding time $\tau_{\mathrm{h}}$ Hanson et al. (2003); Banszerus et al. (2022); Gächter et al. (2022). We convert the current $I$ into the number of charge carriers tunneling through the QD per pulse cycle, $\langle n\rangle/\text{pulse}=I(\tau_{\text{i}}+\tau_{\text{h}}+\tau_{\text{m}})/e$. The number of charge carriers $\langle n\rangle_{\ket{K\uparrow}_{\mathrm{m}}}$ tunneling via the ES |$\ket{K\uparrow}_{\mathrm{m}}$, is directly proportional to the probability of |$\ket{K\uparrow}$ remaining occupied after $\tau_{\text{h}}$, $P_{\ket{K\uparrow}}(\tau_{\text{h}})$. The relative occupation probability of |$\ket{K\uparrow}$ as a function of $\tau_{\mathrm{h}}$ decays exponentially with the characteristic relaxation time, $T_{1}$ Hanson et al. (2003); Banszerus et al. (2022):
$\frac{\langle n\rangle_{\ket{K\uparrow}_{\mathrm{m}}}(\tau_{\mathrm{h}})}{\langle n\rangle_{\ket{K\uparrow}_{\mathrm{m}}}(0)}=\frac{P_{\ket{K\uparrow}}(\tau_{\text{h}})}{P_{\ket{K\uparrow}}(0)}=e^{-\tau_{\mathrm{h}}/T_{1}}.$ (1)

### v2

@@ Abstract
The valley degree of freedom in 2D semiconductors, such as gapped bilayer graphene (BLG) and transition metal dichalcogenides, is a promising carrier of quantum information in the emerging field of valleytronics. While valley dynamics have been extensively studied for moderate band gap 2D semiconductors using optical spectroscopy techniques, very little is known about valley lifetimes in narrow band gap BLG, which is difficult to study using optical techniques. Here, we report single-particle valley relaxation times ($T_{1}$) exceeding several microseconds in electrostatically defined BLG quantum dots (QDs) using a pulse-gating technique. The observed dependence of $T_{1}$ on perpendicular magnetic field can be understood qualitatively and quantitatively by a model in which $T_{1}$ is limited by electron-phonon coupling. We identify the coupling to acoustic phonons via the bond length change and via the deformation potential as the limiting mechanisms.
Charge carriers in 2D materials with a hexagonal crystal lattice have, in addition to the spin, a tunable valley degree of freedom. This renders these materials promising candidates for valleytronics [1, 2, 3, 4], where the valley relaxation time ($T_{1}$) is an important figure of merit, allowing to assess the potential for valley-based information storage. In 2D transition metal dichalcogenides with moderate band gaps, the valley degree of freedom is well accessible for optical manipulation and readout. This allows pump-probe spectroscopy experiments [5, 6], in particular time-resolved Kerr rotation [7, 8, 9, 10, 11], which revealed relaxation times ranging from nanoseconds to microseconds depending on the material and excitation conditions. However, such optical techniques are not readily applicable to narrow-gap BLG, resulting in a lack of knowledge about valley lifetimes in BLG. Recent advances in the confinement of carriers in BLG using quantum dots (QDs), however, are opening up new avenues for the study of valley lifetimes in BLG.
@@
Bernal-stacked BLG comes as a gapless semimetal, where electrons and holes can be described as massive Dirac fermions [12]. However, when applying an out-of-plane electric displacement field $D$, the inversion symmetry of the crystal lattice is broken, as the on-site energy of carbon atoms of the top layer becomes different from that of the atoms of the bottom layer (see Fig. 1(a)) [13]. This leads to the opening of a band gap at the two valleys, $K$ and $K^{\prime}$ (see Fig. 1(b)), which depends on the strength of the symmetry breaking potential, i.e., on $D$  [14, 15, 16], resulting in a tunable band structure that allows for gate-defined charge carrier confinement [17, 18, 19]. The broken inversion symmetry also leads to a finite Berry curvature, $\Omega$, near the K points, where $\Omega$ has opposite signs at the $K$ and $K^{\prime}$ points and has mirror symmetry for electrons and holes [13, 20] (see Fig. 1(c)). The Berry curvature gives rise to a valley-dependent anomalous velocity term leading to the valley-Hall effect in bulk BLG [21, 22] and to finite out-of-plane magnetic moments in BLG QDs. These topological orbital magnetic moments, which have opposite signs for $K$ and $K^{\prime}$, couple to an external out-of-plane magnetic field and are the origin of the valley Zeeman effect in BLG QDs [23, 24].
@@
Figure 1: (a) Lattice structure of BLG highlighting the top and bottom layer and a symmetry breaking displacement field, $D$. The electrostatic confinement potential $U$ allows to form a QD. (b) Band structure of gapped BLG. Close to the $K$ and the $K^{\prime}$ valley a band gap opens. (c) The broken inversion symmetry results in a finite Berry curvature, $\Omega$, near the $K$ and $K^{\prime}$ points, which has opposite sign for the two valleys and for electrons and holes. (d) Single-particle spectrum of a BLG QD. At $B_{\perp}=0$, $\Delta_{\text{SO}}$ leads to the formation of two Kramers’ doublets. A finite $B_{\perp}$ results in a spin and valley Zeeman effect, leading to additional energy splittings of $\Delta E_{\text{s}}=g_{\mathrm{s}}\mu_{\mathrm{B}}B$ and $\Delta E_{\text{v}}=g_{\mathrm{v}}\mu_{\mathrm{B}}B$. The arrows depict the transition energies between the ground state |$\ket{K^{\prime}\uparrow}$ and the three excited states.
@@
To create gate-defined QDs in BLG, the electronic wave function needs to be confined by a potential $U(\textbf{r})$ in real space (see Fig. 1(a)) and will be distributed near the $K$ and $K^{\prime}$ points, in k-space. A single-electron or single-hole QD can then be described by the Hamiltonian $H_{\rm QD}=H_{\rm BLG}+H_{\rm Z}+H_{\rm SO}+U({\bf r})$ [25]. Here, $H_{\rm BLG}$ represents the effective $4\times 4$ Hamiltonian of bulk BLG near the $K$- and $K^{\prime}$-points based on the sublattice and layer degrees of freedom and includes the bulk valley Zeeman effect (see Supplementary Material), which will be further modified by the confinement $U({\bf r})$. $H_{\rm Z}$ denotes the spin Zeeman coupling. $H_{\rm SO}$ describes the intrinsic Kane-Mele spin-orbit (SO) coupling, which lifts the zero $B$-field degeneracy of the four single-particle states, leading to the formation of two Kramers’ pairs with opposite spin and valley quantum numbers, (|$\ket{K\uparrow}$,|$\ket{K^{\prime}\downarrow}$) and (|$\ket{K^{\prime}\uparrow}$, |$\ket{K\downarrow}$), separated by the SO gap, $\Delta_{\mathrm{SO}}$. In BLG devices $\Delta_{\mathrm{SO}}$ has typically values in the range of $40-80~{}\,\mu$eV [18, 23, 20, 26, 27]. In Fig. 1(d) we show the single-particle spectrum of a BLG hole QD as function of the out-of-plane magnetic field, $B_{\perp}$. As $B_{\perp}$ couples to both the spin and the valley magnetic moments, we observe linear energy shifts given by $E(B_{\perp})-E(0)=(\pm g_{\mathrm{s}}\pm g_{\mathrm{v}})\mu_{\mathrm{B}}B_{\perp}/2$ [28]. Here, $\mu_{\text{B}}$ is the Bohr magneton, $g_{\mathrm{s}}\approx 2$ is the spin g-factor and the valley g-factor, $g_{\mathrm{v}}$, quantifies the strength of the Berry-curvature induced valley magnetic moment, which can be tuned by the confinement potential of the QD in a range typically between $g_{\text{v}}\approx 10$ and 70 [24, 29, 30]. All this makes the valley degree of freedom in BLG QDs highly tunable, in stark contrast to its behavior in Si, Ge, or SiGe QDs, enabling significant valley polarization at relatively low $B_{\perp}$-fields, as evident by the separation between the |$\ket{K^{\prime}}$ and excited |$\ket{K}$ states shown in Fig. 1(d).
@@
Figure 2: (a) False-colored scanning electron microscopy image of the gate structure of the device. The plunger gate (PG, red) is connected to a bias tee for applying AC and DC signals to the same gate. (b) Current through the device as function of $V_{\text{PG}}$ at a source-drain bias voltage of $V_{\mathrm{SD}}=200~{}\mu$V. The n-type channel is pinched off close to $V_{\text{PG}}=-6.8V$. Upon further decreasing $V_{\text{PG}}$, a hole QD is formed, and Coulomb resonances appear when additional holes are added to the QD (see labels). (c) Top: Schematic of the square pulse applied to the PG characterized by the voltages $V_{\text{i}}$ and $V_{\text{m}}$ and the times $\tau_{\text{i}}$ and $\tau_{\text{m}}$. Bottom: Schematic of the QD states relative to the electrochemical potentials of the leads. (d) Excited state (ES) spectroscopy using transient current measurements. The average number of charge carriers $\langle n\rangle$ tunneling through the QD per pulse is plotted as function of $\Delta V_{\mathrm{PG}}$ for different $V_{\mathrm{AC}}$ ($f=5~{}$MHz, $B_{\perp}=300$ mT). Traces are offset for clarity. |$\ket{K^{\prime}\uparrow}$ denotes current via the ground state. Orange and yellow dashed lines highlight transient currents via excited states (|$\ket{K^{\prime}\downarrow}$, |$\ket{K\uparrow}$ and |$\ket{K\downarrow}$). (e) Energy $\Delta E$ of the ES relative to the GS as a function of $B_{\perp}$. Fitting $\Delta_{\text{SO}}/2+(g_{\mathrm{v}}+g_{\mathrm{s}}/2)\mu_{\mathrm{B}}B$ (orange dashed line) yields $g_{\mathrm{v}}$. The solid lines indicate the energies of the valley ESs deduced from the fits. The inset shows a close-up of the low-$B_{\perp}$ regime.
@@
The fabricated device consists of a flake of BLG encapsulated by two crystals of hexagonal boron nitride (hBN) and placed on a graphite flake acting as a back gate. On top of the van-der-Waals heterostructure, split gates (SGs) are used to gap out the BLG underneath, resulting in a narrow n-type conductive channel connecting the source (S) and drain (D) leads (see Fig. 2(a)). To confine single charge carriers, the band edge profile along the channel can be adjusted using two layers of interdigitated finger gates (FGs) [19, 17]. One of the FGs is used as a plunger gate (PG) to tune the QD, locally overcompensating for the channel potential set by the back gate. The width of the PG measures about 70 nm and the separation of the SGs is around 80 nm, setting an upper limit of the QD radius $r$ to around 30 to 40 nm. The DC potential applied to the plunger gate, $V_{\text{PG}}$, allows to control the charge carrier occupation down to the last hole (see Fig. 2(b)). To study transient transport through the QD, an AC potential, $V_{\mathrm{AC}}$, can be applied to the PG via a bias tee (see Fig. 2(a)). To maximize the transient currents and to study the relaxation dynamics of the QD states, the FGs adjacent to the PG (yellow and blue) are used to reduce the tunnel coupling between the QD and the left and right reservoir, $\Gamma_{\mathrm{L}}$ and $\Gamma_{\mathrm{R}}$ [31, 32, 33, 34]. Details on the device fabrication and the experimental setup are given in the Supplementary Material.
@@
To study the relaxation dynamics of an excited valley state, we first investigate the single-particle spectrum of the QD. For that purpose, we perform excited state transient current spectroscopy measurements by applying a square pulse with a frequency, $f$, (duty cycle 50%) to the PG (see Fig. 2(c)). Fig. 2(d) shows the average number of charge carriers tunneling through the QD per pulse cycle, $\langle n\rangle/\text{pulse}=I/(fe)$, with the elementary charge $e$, as a function of the pulse amplitude, $V_{\mathrm{AC}}$, and the DC plunger gate voltage, $\Delta V_{\mathrm{PG}}$, relative to the Coulomb peak position at $V_{\mathrm{AC}}=0$. At finite $V_{\mathrm{AC}}$, transport via the ground state (GS) may occur when the GS resides within the bias window during either part of the square pulse ($\tau_{\mathrm{i}},\tau_{\mathrm{m}}$), resulting in a splitting of the GS Coulomb peak (|$\ket{K\uparrow}_{i}$ and |$\ket{K\uparrow}_{m}$[35, 32]. Once $V_{\mathrm{AC}}$ becomes large enough such that an excited state (ES) enters the bias window, a transient current via the ES contributes to the overall current and shows up as a resonance in Fig. 2(d) (see dashed lines). From the positions of the two prominent ES resonances we can extract their energies. Fig. 2(e) depicts the energy difference between the ground state (|$\ket{K^{\prime}\uparrow}$) and the first spin ES (yellow data points, |$\ket{K^{\prime}\downarrow}$) and the valley ESs, (red data points |$\ket{K\downarrow}$, |$\ket{K\uparrow}$), determined from measurements as in Fig. 2(d), as a function of $B_{\perp}$. The energy splitting of the spin ES and the GS increases linearly with $B_{\perp}$ due to the spin Zeeman effect. A fit to the data yields a spin g-factor of $g_{\text{s}}=2.0\pm 0.2$ and a zero-field splitting of $\Delta_{\text{SO}}=75~{}\mu$eV, in agreement with the slightly proximity-enhanced Kane-Mele SO coupling [23, 20, 26]. Due to the finite peak width, the energy of the two valley ESs cannot be determined independently. Thus, the data has been fit considering the average energy splitting with a slope corresponding to $g_{\mathrm{v}}+g_{\mathrm{s}}/2$. A valley g-factor of $g_{\mathrm{v}}=30.2\pm 0.2$ has been determined, similar to values reported in earlier works [23, 24].
@@
Figure 3: (a) Top: Schematic of the three-level pulse scheme applied to the PG which is characterized by the voltages $V_{\text{i}},V_{\text{h}}$, $V_{\text{m}}$ and the times $\tau_{\text{i}},\tau_{\text{h}}$ and $\tau_{\text{m}}$. Bottom: Schematic of the QD states relative to the electrochemical potentials in the leads (see text for details). (b) Current $I$ as a function of $V_{\text{PG}}$ while the pulse sequence in a is applied ($B_{\perp}=0.22$ T, $V_{\mathrm{SD}}=10~{}\mu$V). The valley $T_{1}$ time is derived from the amplitude of |$\ket{K\uparrow}_{\text{m}}$. Dashed curves are Lorentzian fits to the peaks. (c) Relative occupation probability of |$\ket{K\uparrow}$ after the holding pulse, $P_{\ket{K\uparrow}}(\tau_{\text{h}})/P_{\ket{K\uparrow}}(0)$ as a function of the holding time, $\tau_{\text{h}}$. The traces are offset for clarity.
@@
Next, to investigate the relaxation dynamics of the observed single-particle valley ES, we apply a three-level pulse scheme to the PG (see Fig. 3(a)) and measure the tunneling current through the QD [33]. During $\tau_{\mathrm{i}}$, the QD is initialized in the empty state. Subsequently, during $\tau_{\mathrm{h}}$, the GS |$\ket{K^{\prime}\uparrow}$, the spin ES |$\ket{K^{\prime}\downarrow}$ and the two valley ESs |$\ket{K\uparrow}$ and |$\ket{K\downarrow}$ are tuned below the electrochemical potentials of the source and drain leads. After the characteristic tunneling time, on the order of $1/(\Gamma_{\mathrm{L}}+\Gamma_{\mathrm{R}})$ either the GS or one of the three ESs will be occupied by a single charge carrier. A charge carrier in an ES has the chance to relax into an energetically lower lying state by either spin or valley relaxation with a characteristic relaxation time, $T_{1}$. Finally, during $\tau_{\mathrm{m}}$, we perform a valley selective readout, measured by aligning the |$\ket{K}$ states in the bias window. Only charge carriers occupying one of the two |$\ket{K}$ states, which have not relaxed into a |$\ket{K^{\prime}}$, can tunnel out and contribute to the transient current. Fig. 3(b) shows the current, $I$, through the QD as a function of $V_{\mathrm{PG}}$ while applying the pulse sequence depicted in Fig. 3(a). The three peaks labeled |$\ket{K^{\prime}\uparrow}_{\mathrm{i}}$, |$\ket{K^{\prime}\uparrow}_{\mathrm{h}}$ and |$\ket{K^{\prime}\uparrow}_{\mathrm{m}}$ correspond to GS transport during each of the three pulse steps. Furthermore, transient currents via the three ESs, |$\ket{K\uparrow}_{\mathrm{m}}$, |$\ket{K\downarrow}_{\mathrm{m}}$ and |$\ket{K^{\prime}\downarrow}_{\mathrm{m}}$, can be observed during $\tau_{\mathrm{m}}$. The relaxation time $T_{1}$ of |$\ket{K\uparrow}$ into an energetically lower lying state can be probed by measuring the amplitude of the |$\ket{K\uparrow}_{\mathrm{m}}$ peak as a function of the holding time $\tau_{\mathrm{h}}$ [31, 33, 34]. We convert the current $I$ into the number of charge carriers tunneling through the QD per pulse cycle, $\langle n\rangle/\text{pulse}=I(\tau_{\text{i}}+\tau_{\text{h}}+\tau_{\text{m}})/e$. The number of charge carriers $\langle n\rangle_{\ket{K\uparrow}_{\mathrm{m}}}$ tunneling via the ES |$\ket{K\uparrow}_{\mathrm{m}}$, is directly proportional to the probability of |$\ket{K\uparrow}$ remaining occupied after $\tau_{\text{h}}$, $P_{\ket{K\uparrow}}(\tau_{\text{h}})$. The relative occupation probability of |$\ket{K\uparrow}$ as a function of $\tau_{\mathrm{h}}$ decays exponentially with the characteristic relaxation time, $T_{1}$, $\langle n\rangle_{\ket{K\uparrow}_{\mathrm{m}}}(\tau_{\mathrm{h}})/\langle n\rangle_{\ket{K\uparrow}_{\mathrm{m}}}(0)=P_{\ket{K\uparrow}}(\tau_{\text{h}})/P_{\ket{K\uparrow}}(0)=e^{-\tau_{\mathrm{h}}/T_{1}}$ [31, 33].